# Query-focused and Memory-aware Reranker for Long Context Processing

### v1

@@ Abstract
Built upon the existing analysis of retrieval heads in large language models, we propose an alternative reranking framework that trains models to estimate passage–query relevance using the attention scores of selected heads. This approach provides a listwise solution that leverages the holistic information within the entire candidate shortlist during ranking. At the same time, it naturally produces continuous relevance scores, enabling training on arbitrary retrieval datasets without requiring Likert-scale supervision. Our framework is lightweight and effective, requiring only small-scale models (e.g., 4B parameters) to achieve strong performance. Extensive experiments demonstrate that our method outperforms existing state-of-the-art pointwise and listwise rerankers across multiple domains, including Wikipedia and long narrative datasets. It further establishes a new state-of-the-art on the LoCoMo benchmark that assesses the capabilities of dialogue understanding and memory usage. We further demonstrate that our framework supports flexible extensions. For example, augmenting candidate passages with contextual information further improves ranking accuracy, while training attention heads from middle layers enhances efficiency without sacrificing performance [^1] The models are available at https://huggingface.co/MindscapeRAG/QRRanker.
@@ 1 Introduction
A long line research applies an additional reranker module on the shortlist returned from embedding models to resolve this challenge. The rerankers use larger models, more powerful representations (like cross-attention). The fast development of LLMs boosts many LLM-based reranker releases to benefit from the reasoning capabilities of LLMs Zhang et al. (2025b); Sun et al. (2025); Liu et al. (2025a); Pradeep et al. (2023b). These rerankers can adopt either pointwise or listwise formulations. Pointwise lost the global view of the shortlist, but can give scores. Listwise approaches, on the other hand, directly inherit the long-context reasoning and text generation ability of the backbone LLMs, which takes a holistic view of the shortlist, but the next-token prediction limits the prediction of fine-grained scores, and the predicted float numbers cannot always accurately reflect the true confidence Liu et al. (2025b); Lin et al. (2024). As a result, they adopt a Likert rating regime, asking the models to output a five-point or ten-point scale score for each input document. Which limited the available training data.
@@
Experiments on various domains, including Wikipedia QA tasks (Musique Trivedi et al. (2022), HotpotQA Yang et al. (2018)), long narrative QA tasks (NarrativeQA Kočiskỳ et al. (2018), DetectiveQA Xu et al. (2025b)) and long-context dialogue (LoCoMo Maharana et al. (2024)), demonstrate the advantage of our QRRanker. As a versatile ranking framework, our approach not only outperforms the state-of-the-art general-purpose pointwise and listwise models like Qwen-Rerank and GroupRank, but also consistently improves over the domain-specific ranking approaches, such as HippoRAG-v2 Guti’errez et al. (2025) for Wikipedia QA and a list of recent memory-enhanced approaches Li et al. (2025b); Rasmussen et al. (2025); Hu et al. (2026a) on LoCoMo.
@@ 2 Related Work
Ranking techniques are acceptedly constructed based on two structures: Siamese network (Bi-encoder; Koch et al. 2015) and Cross-encoder (Thakur et al., 2021). Embedding models (Zhang et al., 2025b) is the first one usually used to rank the whole corpus of documents with embeddings stored for reuse. However, they are limited by the “geometric bottleneck”, failing to encode more fine-grained interactions between query and document. The limitation can be alleviated by cross-encoders, which score every document by cross-attention with the query. The computing burden of dedicated re-encoding every pair of query-document narrows the way for cross-encoders only reranking top-n documents ranked by bi-encoders, which produces refined sorting of top docs. Therefore, cross-encoders are called Rerankers.
@@
Memory construction and utilization to alleviate problems of long-context processing become a hot spot nowadays. For long story understanding, Li et al. (2025a) construct global memory to enhance retrieval and generation. For dialogue management, sophisticated graphs (Jiang et al., 2026; Xu et al., 2025a; Rasmussen et al., 2025; Hu et al., 2026b, a), trees (Li et al., 2026), and systems Chhikara et al. (2025); Li et al. (2025b); Nan et al. (2025); Tao et al. (2026); Zou et al. (2026) of events, personas, and chunks are designed to accurately extract related dialogue history for further use. However, a better and powerful search for history, with simple memory construction, can beat complicated memory management, and we will show our solution to reach this goal.
@@ 3 Preliminaries: QR-head
QR heads compute the retrieval score for a chunk $c_{i}$ in a similar way like Eq. 1 by replacing $\sum_{c\in G}$ with $\sum_{h\in H_{QR}}$. Zhang et al. (2025a) further add a score calibration to mitigate intrinsic biases in attention weights, which encodes a null query NN=“N/A” and subtracts its $\frac{1}{|N|}\sum_{w_{q}\in N}A^{N\rightarrow c_{i}}$. Notably, with our QR training, calibration becomes optional.
@@ 4.1 Data Construction for QR Training
2:Training set $\mathcal{T}$
[…]
4:for all question $Q$ in $\mathcal{D}$ do
6:  $C\leftarrow\mathcal{R}(Q,K)$ $\triangleright$ retrieve top-$K$ candidate chunks
7:  for all $c_{i}\in C$ do
8:   $y_{i}\leftarrow\mathbb{I}[c_{i}\in G]$
9:  end for
10:  if $\mathsf{UseMem}$ then
13:  else
14:   $M\leftarrow\emptyset$
15:  end if
16:  $\mathcal{T}\leftarrow\mathcal{T}\cup\{(Q,M,C,\{y_{i}\}_{i=1}^{K})\}$
17:end for
18:return $\mathcal{T}$
@@ 4.1.1 Listwise Training Instances
Optionally, we construct a summary prefix by mapping the retrieved chunks to their corresponding summaries, and prepend these summaries before the chunk list, i.e., $X=[M;C]$. Alg. 1 summarizes this construction on NarrativeQA; MuSiQue follows the same procedure except that relevant evidence directly comes from their official supporting facts. We describe how the summaries are constructed in the next subsection.
@@ 4.1.2 Summary Construction
For long narrative books, we construct block-level summaries that respect the sequential nature of storytelling. Specifically, we split each book into blocks (20 consecutive chunks per block) and generate one summary per block. (see Appendix A.1)
@@
For dialogue-based data, we extract structured events from conversations and form an event-centric summary. Each event is represented by a short description and is linked to its source utterances, enabling traceability to the original dialogue. (see Appendix A.2).
@@ 4.2 QR Training
Obtaining QR heads precomputed by the QR score mentioned in Sec. 3, our training scheme focuses on training these heads. For a question $Q$ and the top 50 candidate documents $C=[c_{1},...,c_{50}]$ ranked by a retriever (e.g., embedding models like Qwen3-Embedding), where gold (positive) documents are $G=[c_{g0},..,c_{gm}]$, the prompt input to QRRanker is constructed by concatenating $C$ and $Q$ in order with some instructions: $\texttt{P}=\text{Inst}(C,Q)$, where the instruction template is provided in Appendix A.3.
@@ 5.1 Datasets
Long-context dialogue memory
We evaluate our model on LoCoMo Maharana et al. (2024), a large-scale benchmark designed for long-context dialogue memory. The dataset comprises 50 multi-session dialogues across 10 distinct user groups, with each dialogue averaging approximately 9,000 tokens. Following prior work, we report performance across four fine-grained categories: single-hop, multi-hop, temporal reasoning, and open-domain.
@@ 5.2 Baselines
For the long-context dialogue task on LoCoMo, we compare QRRanker with a range of strong baselines, including: A-Mem Xu et al. (2025a), MemoryOS Li et al. (2025b), Zep Rasmussen et al. (2025), Mem0 Chhikara et al. (2025), Nemori Nan et al. (2025), and LightMem Fang et al. (2025); TiMem Li et al. (2026), Synapse Jiang et al. (2026), Membox Tao et al. (2026), CompassMem Hu et al. (2026b), and ES-Mem Zou et al. (2026); SimpleMem Liu et al. (2026). Detailed baseline descriptions are provided in Appendix B.
@@ 5.3 Implementation Details
Our QRRanker is trained on Qwen3-4B-Instruct-2507, with QR heads selected as described in Appendix C. In the training process, the scalescale factor in the max-min norm is set to 8; the batch size is set to 1; the gradient accumulating step is set to 4; the learning rate is set to 1e-5. We utilize the DeepSpeed ZERO2 strategy and train QRRanker using 8 H20 GPUs.
For downstream QA evaluation, we use task-specific prompting for generation; the full prompt templates for NarrativeQA, DetectiveQA, and LoCoMo are provided in Appendix A. We employ Qwen3-8B as the generator for NarrativeQA and DetectiveQA, where books are chunked into non-overlapping passages of $\sim$200 tokens. For the LoCoMo benchmark, we utilize GPT-4o-mini and GPT-5-mini as the generators. We segment the dialogue history into small chunks, ensuring that utterance continuity is preserved, with an average chunk size of 258 tokens. When enabling the memory-aware setting, we prepend a summary prefix before the ranked chunk list. We cap the summary prefix at 512 tokens and select summaries based on their coverage of the retrieved/reranked chunks.
@@ 6.1 Main Results
We conduct extensive experiments spanning three distinct domains: Wikipedia multi-hop QA, long-context story QA, and dialogue memory. These experiments cover five datasets in both English and Chinese. Tables 1 and 2 present the overall reranking performance measured by Recall@$k$. The results demonstrate that our proposed QRRanker consistently achieves the best performance across all datasets. It substantially outperforms embedding-only retrieval, strong reranking baselines such as Qwen-Reranker, and the vanilla out-of-box QRHeads variant. Furthermore, we evaluate downstream generation on narrative QA and dialogue memory as reported in Tables 3 and 4. In these tasks, our QRRanker yields consistent gains and demonstrates strong generalization from retrieval reranking to end tasks
@@
We first analyze the retrieval effectiveness of QRRanker when applied to rerank the top-50 candidates retrieved by embeddings. As shown in Table 1, QRRanker establishes a new state-of-the-art benchmark. It surpasses the strong baseline Qwen-Reranker-4B by a substantial margin and improves the average recall significantly. On Wikipedia datasets such as Musique and HotpotQA, QRRanker outperforms complex graph-based methods like HippoRAG Guti’errez et al. (2025). Remarkably, it also exceeds the performance of GroupRank-32B despite being significantly more lightweight. This indicates that our method captures inter-passage dependencies more effectively than simple groupwise scoring or graph traversal. The performance gap is particularly evident in the Story domain, where context tracking is critical. For instance, QRRanker achieves a Recall@10 of 54.93 on NarrativeQA compared to 48.83 for GroupRank and 48.89 for the vanilla QRHeads. Finally, on LoCoMo (Table 2), QRRanker maintains the same advantage, indicating its effectiveness in retrieving relevant context from long conversational histories.
@@
As summarized in Table 3, QRRanker demonstrates superior efficiency on LoCoMo, achieving the best Overall F1 with a highly compact input budget. By using only 854 tokens on average (top-3 chunks) directly from the raw dialogue history, our approach achieves an Overall F1 of 57.03 with GPT-4o-mini and 57.32 with GPT-5-mini. In contrast, many memory-augmented frameworks require substantially larger budgets to maintain explicit memory stores or graphs. Our approach instead reranks the top-50 chunks retrieved by the embedding retriever and feeds only a small set of top-ranked raw dialogue chunks to the generator. This lightweight design preserves high inference efficiency and low system complexity while still capturing long-range dependencies, yielding the highest Overall F1 among prior reported results on LoCoMo in our comparison.
@@ 6.2 Results with Contextual Information
Table 5: Recall@3 comparison of QRRanker with chunk-only inputs versus a summary prefix (+Sum) as contextual memory. $\Delta$ indicates the absolute change after adding the summary prefix.
@@ 6.3 Results with Heads from Different Layer- Levels
QRRanker uses static preset heads, which invokes our curiosity about the heads from which level of layers are suitable as starters for QR training. We propose a variant that dynamically selects heads from a range of continuous layers for every sample. The variant totally picks up 16 heads from layer $l_{s}$ to $l_{e}$ with $16/(l_{e}-l_{s})$ heads per layer, where $l_{s}$-$l_{e}$ determines the level of layers (i.e., low, middle, high). Details of the variant are elaborated in Appendix D. We train and evaluate both QRRanker and its variants on the NarrativeQA dataset.
@@
As shown in Tab. 6, training models with lower layers 10-17 shows a significant performance drop, while middle layers 17-24 and top layer 28-35 almost keep the same performance as QRRanker. Intuitively, lower layers truncate too much knowledge from higher layers, and heads in the middle-to-top layers are more likely to be retrievers. The outcome aligns with the phenomenon that QR heads in QRRanker are all positioned in the middle layers (17-24). Interestingly, we compare QR heads in QRRanker with those selected by the variant (17-24), and the degree of overlap is pretty low. It indicates that, with QR training, such potential is activated, which shows that our method can utilize the robustness of heads, even not QR heads, from the middle to the top. This provides a way to only focus on heads in the middle and truncate the higher layers for a smaller and faster ranker. We quantify the inference efficiency benefits of this middle-layer truncation in Section 6.4.
@@ 6.4 Inference Efficiency
We further investigate the computational efficiency of our approach compared to baselines on a set of 20 queries. As shown in Table 7, QRRanker achieves lower P50/P95 latency than Qwen3-Reranker-4B, while also reducing compute (TFLOPs) and peak memory. Moreover, QRRanker(middle) further improves efficiency by truncating the model after layer 24, discarding higher layers. It achieves the best P50/P95 latency with additional reductions in compute and memory. For Qwen3-Reranker-4B, we report two inference settings. With batch=50, all 50 chunk–query pairs are processed in a single forward pass. With batch=1, the 50 pairs are processed with 50 separate forward passes, which substantially increases latency. Overall, QRRanker provides a better performance and cost trade-off, and the truncated middle-layer variant offers an especially lightweight and fast option.
@@
Table 7: Inference efficiency comparison in latency (P50/P95), compute (TFLOPs per query), and peak GPU memory. All models are evaluated under the same hardware and inference settings over 20 queries. For Qwen3-Reranker-4B, batch=50 processes 50 chunk–query pairs in a single forward pass, whereas batch=1 processes the 50 pairs with 50 separate forward passes. QRRanker(middle) truncates the model after layer 24.

### v2

@@ Abstract
Built upon the existing analysis of retrieval heads in large language models, we propose an alternative reranking framework that trains models to estimate passage–query relevance using the attention scores of selected heads. This approach provides a listwise solution that leverages the holistic information within the entire candidate shortlist during ranking. At the same time, it naturally produces continuous relevance scores, enabling training on arbitrary retrieval datasets without requiring Likert-scale supervision. Our framework is lightweight and effective, requiring only small-scale models (e.g., 4B parameters) to achieve strong performance. Extensive experiments demonstrate that our method outperforms existing state-of-the-art pointwise and listwise rerankers across multiple domains, including Wikipedia and long narrative datasets. It further establishes a new state-of-the-art on the LoCoMo benchmark that assesses the capabilities of dialogue understanding and memory usage. We further demonstrate that our framework supports flexible extensions. For example, augmenting candidate passages with contextual information further improves ranking accuracy, while training attention heads from middle layers enhances efficiency without sacrificing performance.
@@ 1 Introduction
A long line research applies an additional reranker module on the shortlist returned from embedding models to resolve this challenge. The rerankers use larger models, more powerful representations (like cross-attention). The fast development of LLMs boosts many LLM-based reranker releases to benefit from the reasoning capabilities of LLMs Zhang et al. (2025b); Sun et al. (2025); Liu et al. (2025a); Pradeep et al. (2023b). These rerankers can adopt either pointwise or listwise formulations. Pointwise lost the global view of the shortlist, but can give scores. Listwise approaches, on the other hand, directly inherit the long-context reasoning and text generation ability of the backbone LLMs, which takes a holistic view of the shortlist, but the next-token prediction limits the prediction of fine-grained scores, and the predicted float numbers cannot always accurately reflect the true confidence Liu et al. (2025b); Lin et al. (2024). As a result, they adopt a Likert rating regime, asking the models to output a five-point or ten-point scale score for each input document, which limits the available training data.
@@
Experiments on various domains, including Wikipedia QA tasks (Musique Trivedi et al. (2022), HotpotQA Yang et al. (2018)), long narrative QA tasks (NarrativeQA Kočiskỳ et al. (2018), DetectiveQA Xu et al. (2025b)) and long-term dialogue (LoCoMo Maharana et al. (2024)), demonstrate the advantage of our QRRanker. As a versatile ranking framework, our approach not only outperforms the state-of-the-art general-purpose pointwise and listwise models like Qwen-Rerank and GroupRank, but also consistently improves over the domain-specific ranking approaches, such as HippoRAG-v2 Guti’errez et al. (2025) for Wikipedia QA and a list of recent memory-enhanced approaches Li et al. (2025b); Rasmussen et al. (2025) on LoCoMo.
@@ 2 Related Work
Ranking techniques are commonly constructed based on two structures: Siamese network (Bi-encoder; Koch et al. 2015) and Cross-encoder (Thakur et al., 2021). Embedding models (Zhang et al., 2025b) is the first one usually used to rank the whole corpus of documents with embeddings stored for reuse. However, they are limited by the “geometric bottleneck”, failing to encode more fine-grained interactions between query and document. The limitation can be alleviated by cross-encoders, which score every document by cross-attention with the query. The computing burden of dedicated re-encoding every pair of query-document narrows the way for cross-encoders only reranking top-n documents ranked by bi-encoders, which produces refined sorting of top docs. Therefore, cross-encoders are called Rerankers.
@@
Memory construction and utilization to alleviate problems of long-context processing has become a hot spot nowadays. For long story understanding (Zhou et al., 2025; Kočiskỳ et al., 2018; Xu et al., 2025b; Wu et al., 2025), Li et al. (2025a) construct global memory to enhance retrieval and generation. For dialogue management, sophisticated graphs (Jiang et al., 2026; Xu et al., 2025a; Rasmussen et al., 2025; Hu et al., 2026b, a; Zhou et al., 2026), trees (Li et al., 2026), and systems Chhikara et al. (2025); Li et al. (2025b); Nan et al. (2025); Tao et al. (2026); Zou et al. (2026) of events, personas, and chunks are designed to accurately extract related dialogue history for further use. However, a more powerful search for history, with simple memory construction, can beat complicated memory management, and we will show our solution to reach this goal.
@@ 3 Preliminaries: QR-head
QR heads compute the retrieval score for a chunk $c_{i}$ in a similar way like Eq. 1 by replacing $\sum_{c\in G}$ with $\sum_{h\in H_{QR}}$. Zhang et al. (2025a) further adds a score calibration to mitigate intrinsic biases in attention weights, which encodes a null query NN=“N/A” and subtracts its $\frac{1}{|N|}\sum_{w_{q}\in N}A^{N\rightarrow c_{i}}$. Notably, with our QR training, calibration becomes optional.
@@ 4.1.1 Listwise Training Instances
Optionally, we build a summary prefix by mapping the retrieved chunks to their corresponding summaries and prepending these summaries to the chunk list, i.e., $X=[M;C]$. The detailed construction procedure for NarrativeQA is provided in alg. 1 in Appendix A. MuSiQue follows the same pipeline, except that relevant evidence is taken directly from its official supporting facts. We describe the summary construction process in the next subsection.
@@ 4.1.2 Summary Construction
For long narrative books, we employ a block-based summarization strategy that respects the temporal flow of the narrative. Specifically, we segment each book into blocks (20 consecutive chunks per block) and generate a corresponding summary for each. This process is detailed in Appendix B.1.
@@
For dialogue-based data, we extract structured events from conversations and form an event-centric summary. Each event is represented by a short description and is linked to its source utterances, enabling traceability to the original dialogue. (Refer to Appendix B.2 for details).
@@ 4.2 QR Training
Obtaining QR heads precomputed by the QR score mentioned in § 3, our training scheme focuses on training these heads. For a question $Q$ and the top 50 candidate documents $C=[c_{1},...,c_{50}]$ ranked by a retriever (e.g., embedding models like Qwen3-Embedding), where gold (positive) documents are $G=[c_{g0},..,c_{gm}]$, the prompt input to QRRanker is constructed by concatenating $C$ and $Q$ in order with some instructions: $\texttt{P}=\text{Inst}(C,Q)$, where the instruction template is provided in Appendix B.3.
@@ 5.1 Datasets
Long-term dialogue memory
We evaluate our model on LoCoMo Maharana et al. (2024), a large-scale benchmark designed for long-term dialogue memory. The dataset comprises 50 multi-session dialogues across 10 distinct user groups, with each dialogue averaging approximately 9,000 tokens. Following prior work, we report performance across four fine-grained categories: single-hop, multi-hop, temporal reasoning, and open-domain.
@@ 5.2 Baselines
For the long-term dialogue task on LoCoMo, we compare QRRanker with a range of strong baselines, including: A-Mem Xu et al. (2025a), MemoryOS Li et al. (2025b), Zep Rasmussen et al. (2025), Mem0 Chhikara et al. (2025), Nemori Nan et al. (2025), and LightMem Fang et al. (2025); TiMem Li et al. (2026), Synapse Jiang et al. (2026), Membox Tao et al. (2026), CompassMem Hu et al. (2026b), and ES-Mem Zou et al. (2026); SimpleMem Liu et al. (2026). Detailed baseline descriptions are provided in Appendix C.
@@ 5.3 Implementation Details
Our QRRanker is trained on Qwen3-4B-Instruct-2507, with QR heads selected as described in Appendix D. In the training process, the scalescale factor in the max-min norm is set to 8; the batch size is set to 1; the gradient accumulating step is set to 4; the learning rate is set to 1e-5. We utilize the DeepSpeed ZERO2 strategy and train QRRanker using 8 H20 GPUs.
For downstream QA evaluation, we use task-specific prompting for generation; the full prompt templates for NarrativeQA, DetectiveQA, and LoCoMo are provided in Appendix B. We employ Qwen3-8B as the generator for NarrativeQA and DetectiveQA, where books are chunked into non-overlapping passages of $\sim$200 tokens. For the LoCoMo benchmark, we utilize GPT-4o-mini and GPT-5-mini as the generators. We segment the dialogue history into small chunks, ensuring that utterance continuity is preserved, with an average chunk size of 258 tokens. When enabling the memory-aware setting, we prepend a summary prefix before the ranked chunk list. We cap the summary prefix at 512 tokens and select summaries based on their coverage of the retrieved/reranked chunks.
@@ 6.1 Main Results
We conduct experiments across three representative long-context settings: multi-hop question answering over Wikipedia, long-story question answering, and dialogue memory, covering five datasets in both English and Chinese. Tables 1 and 2 summarize reranking performance in terms of Recall@$k$, while Tables 3 and 4 report downstream generation results. Overall, QRRanker consistently achieves the best overall results across settings, demonstrating improvements in both retrieval quality and downstream task performance.
@@
We first analyze the retrieval effectiveness of QRRanker when applied to rerank the top-50 candidates retrieved by SFT-Embedding-8B. As shown in Table 1, QRRanker establishes a new state-of-the-art benchmark. It surpasses the Qwen-Reranker-4B by a substantial margin and improves the average recall significantly. On Wikipedia datasets such as Musique and HotpotQA, QRRanker outperforms complex graph-based methods like HippoRAG Guti’errez et al. (2025). Remarkably, it also exceeds the performance of GroupRank-32B Sun et al. (2025) despite being significantly more lightweight. This indicates that our method captures inter-passage dependencies more effectively than simple groupwise scoring or graph traversal.
The performance gap is particularly evident in the story domain, where context tracking is critical. For instance, QRRanker achieves a Recall@10 of 54.93 on NarrativeQA compared to 48.83 for GroupRank and 48.89 for the out-of-box QRHeads. Moreover, Table 2 shows that QRRanker maintains the same advantage on LoCoMo, further demonstrating its effectiveness in retrieving relevant context from long conversational histories.
@@
As summarized in Table 3, QRRanker achieves the best Overall F1 on LoCoMo with a highly compact input budget. Using only 854 tokens on average (top-3 chunks) from the raw dialogue history, it attains 57.03 Overall F1 with GPT-4o-mini and 57.32 with GPT-5-mini. In contrast, many memory-augmented baselines require substantially larger token budgets to maintain explicit memory stores or graph structures. QRRanker instead reranks the top-50 chunks retrieved by the embedding retriever and feeds only a few top-ranked raw dialogue chunks to the generator. This simple and lightweight design remains highly effective at capturing long-range dependencies, yielding the superior overall performance in our LoCoMo comparison.
@@ 6.2 Results with Contextual Information
Table 5: Recall@3 comparison of QRRanker with chunk-only inputs versus a summary prefix (+Sum) as contextual memory.
@@ 6.3 Results with Heads from Different Layer- Levels
QRRanker uses static preset heads, which invokes our curiosity about the heads from which level of layers are suitable as starters for QR training. We propose a variant that dynamically selects heads from a range of continuous layers for every sample. The variant totally picks up 16 heads from layer $l_{s}$ to $l_{e}$ with $16/(l_{e}-l_{s})$ heads per layer, where $l_{s}$-$l_{e}$ determines the level of layers (i.e., low, middle, high). Details of the variant are elaborated in Appendix E. We train and evaluate both QRRanker and its variants on the NarrativeQA dataset.
@@
As shown in Tab. 6, training models with lower layers 10-17 shows a significant performance drop, while middle layers 17-24 and top layer 28-35 almost keep the same performance as QRRanker. Intuitively, lower layers truncate too much knowledge from higher layers, and heads in the middle-to-top layers are more likely to be retrievers. The outcome aligns with the phenomenon that QR heads in QRRanker are all positioned in the middle layers (17-24). Interestingly, we compare QR heads in QRRanker with those selected by the variant (17-24), and the degree of overlap is pretty low. It indicates that, with QR training, such potential is activated, which shows that our method can utilize the robustness of heads, even not QR heads, from the middle to the top. This provides a way to only focus on heads in the middle and truncate the higher layers for a smaller and faster ranker. We quantify the inference efficiency benefits of this middle-layer truncation in § 6.4.
@@ 6.4 Inference Efficiency
We compare the computational efficiency of our method with baselines on 20 random samples. Table 7 shows that QRRanker achieves lower P50/P95 latency than Qwen3-Reranker-4B, while also reducing TFLOPs and peak memory. Moreover, QRRanker(middle) further improves efficiency by truncating the model after layer 24, discarding higher layers. It achieves the best P50/P95 latency with additional reductions in compute and memory. For Qwen3-Reranker-4B, we report two inference settings. With batch=50, all 50 chunk–query pairs are processed in a single forward pass. With batch=1, the 50 pairs are processed with 50 separate forward passes, which substantially increases latency. Overall, QRRanker provides a better performance and cost trade-off, and the truncated middle-layer variant offers an especially lightweight and fast option.
@@
Table 7: Inference efficiency comparison in latency (P50/P95), compute (TFLOPs per query), and peak GPU memory. All models are evaluated under the same hardware and inference settings over 20 queries. QRRanker(middle) truncates the model after layer 24.